OpenStudy (anonymous):

Find the measure of each angle: m<9 = 2x - 4 m<10 = 2x + 4 They are supplementary angles. I'm lost.

OpenStudy (anonymous):

Here's what I got: m<9 + m<10 = 180 m<9 + 2x + 4 = 180 m<9 + 2x = 176 I'm not sure what to do next.

OpenStudy (anonymous):

rewrite m<9 as 2x - 4

OpenStudy (anonymous):

so you'll have 2x - 4 + 2x = 176

OpenStudy (anonymous):

so next step is to solve for x

OpenStudy (anonymous):

wait. let me figure it out

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

where'd you get the + 2x?

OpenStudy (anonymous):

from what you wrote

OpenStudy (anonymous):

m<9 + 2x = 176

OpenStudy (anonymous):

i just rewrote m<9 as 2x - 4

OpenStudy (anonymous):

is it 44?

OpenStudy (anonymous):

are we talking about the value of x?

OpenStudy (anonymous):

2x - 4 + 2x = 176 add 4 to both sides 2x + 2x = 180 combine like terms 4x = 180 divide both sides by 4 x should be 45

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

wait lol i got confused. can we start at the beginning? "_"

OpenStudy (anonymous):

sorry i have to leave

OpenStudy (anonymous):

aight but ur process is right, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

right. k lol

OpenStudy (anonymous):

once you find x just substitute x into 2x - 4 and 2x + 4

OpenStudy (anonymous):

oki thanks